Suppose that the functions q and r are defined as follows q (x)=x2 +7, r(x)=√x+4. Find the following.

(r.q)(5)=

(q.r)(5)=
Answer #1

SOLUTION;

q(x)=x^{2}+7

r(x)=\sqrt{x+4}

Now,

r(5)=\sqrt{5+4}=3 or -3

q(5)=(5)^{2}+7=32

Now,

a)

(r\circ q)(5)=r(q(5))=r(32)=\sqrt{32+4}

(r\circ q)(5)=\sqrt{36}

(r\circ q)(5)= 6. OR. -6

And,

b)

  (q\circ r)(5)= q(r(5))=q(3)= 3^{2}+7=16

Or, (q\circ r)(5)= q(r(5))=q(-3)= (-3)^{2}+7=16 answe is

(q\circ r)(5)=16